OpenStudy (anonymous):

what is the function of ? y=999x-3481 helpp!!?

OpenStudy (loser66):

line

OpenStudy (sweetburger):

much line very positive such increasing

OpenStudy (anonymous):

Tell whether each equation is a function. Use the vertical line test to help you. Equation Is it a function? y=999x-3481 y-0.3x=-4 y-x=sin⁡(x) y=√(16-x^2 )+3 x^2/68+y^2/17=1 x*y=5 this ????